# CSRv2: Unlocking Ultra-Sparse Embeddings

### v1

@@ 1 Introduction
In the era of large foundation models, the quality of embeddings has become a decisive factor shaping downstream performance across tasks such as retrieval, classification and recommendation. Yet the dominant practice still relies on dense representations with thousands of dimensions (e.g., 2048 – 8192). While highly expressive, such embeddings incur substantial costs in storage, memory, and inference latency. These inefficiencies are magnified in large-scale and real-time deployments, where embedding computation and serving often dominate system throughput. As models scale further, embedding efficiency emerges as a central bottleneck – limiting both web-scale applications and deployment on resource-constrained platforms such as mobile and edge devices.
@@ 3.1 Tackling Massive Dead Neurons with k-Annealing
(b) Accuracy vs active dim
(c) Dead neuron during training
@@
We find this approach effectively maintains a low dead-neuron rate during training. As shown in Figure 2(c), although dead neurons rise slightly at target sparsity, their final proportion is far lower than training directly with $k_{\mathrm{final}}$. This indicates that a curriculum schedule provides richer gradients and avoids collapse into the dead-neuron regime. Similar to simulated annealing, a larger $k_{\mathrm{init}}$ promotes exploration and diverse activations, while annealing gradually sharpens embeddings toward the ultra-sparse regime. Figure 2(b) confirms this, as k-annealing yields consistent performance gains across sparsity levels.
@@ C.3 Implementation Details
To ensure a fair comparison between MRL and CSRv2 on the MTEB benchmark—particularly with respect to domain alignment—we select a backbone model that is not natively supported by MRL: e5-Mistral-7B (Wang et al., 2023). We then finetune this model on a carefully curated collection of multi-domain datasets. Specifically, the training data is drawn from three complementary sources: (i) datasets included in the MTEB benchmark (Muennighoff et al., 2022), (ii) the embedding training data collection curated by the Sentence Transformers team [^5] https://huggingface.co/datasets/sentence-transformers/embedding-training-data , and (iii) a suite of public retrieval datasets introduced in Zhang et al. (2025a). During preprocessing, we first deduplicate datasets that appear across multiple collections. Subsequently, following the natural supervision strategy outlined in Section 3.2, we sample up to 20,000 sentence pairs per dataset, resulting in a consolidated training corpus of approximately one million examples.
We finetune e5-Mistral-7B on this corpus using a batch size of 2,048—a scale commonly adopted by existing MRL-compatible models. Full details of the hyperparameter configuration are provided in Table 5. In contrast, the Qwen3-Embedding-4B model (Zhang et al., 2025b) already incorporates native MRL support; thus, no additional finetuning is required for this backbone.
@@ D.1 Evaluation Metrics
Following the methodology established by Kusupati et al. (2022), we adopt 1-nearest neighbor (1-NN) accuracy as the primary metric for evaluating visual representations. This metric is computed using FAISS (Jacob et al., 2018) with exact L2 distance search. In contrast to classification accuracy—which depends on the specific architecture and training procedure of a downstream classifier—1-NN accuracy provides a direct assessment of whether semantically similar instances are embedded in close proximity within the representation space. Consequently, it serves as a more model-agnostic and training-free probe of intrinsic representation quality.
@@ G.3 Analysis on Unbalanced Weightable Settings
We propose MRL-reweight, where we follow MRL’s settings and set the following weights $\{5,4,3,2,1,\ldots,1\}$ for dimensions 2, 4, 8, 16, 32, and so on, up to 4096. As shown in Table 17, applying larger weights at earlier stages does, to some extent, improve the performance of MRL on low-dimensional scales. However, while MRL-reweight offers some improvements, the performance does not quite reach the level of CSRv2. We hypothesize that this discrepancy arises because sparse vectors—being more comprehensive in capturing feature combinations—are more difficult to achieve with a truncated representation (e.g., retaining only the first 8 values).
@@ H.3 Potential Applications of CSRv2 in Vector Quantization
Vector quantization (VQ) methods—including Product Quantization (PQ) (Jegou et al., 2010), Optimized Product Quantization (Ge et al., 2013), and more recent anisotropic schemes such as AVQ (Guo et al., 2020), are central to real-world large-scale vector search systems where memory footprint, latency, and hardware efficiency are critical. While our main work focuses on the role of ultra-sparse representations in improving retrieval quality and compute efficiency, it is worth noting that CSRv2 is highly compatible with these widely-used quantization techniques.
CSRv2’s ultra-sparse structure—activating only $k\in\{2,4,8\}$ dimensions out of a large latent space—naturally complements vector quantization methods such as PQ and AVQ. As only a few coordinates are non-zero, quantization can be applied exclusively to these active values (or their indices), enabling a two-stage compression pipeline of sparsity + quantization that substantially reduces both memory and lookup cost. Unlike dense embeddings (e.g., MRL), where quantization error spreads across all dimensions, CSRv2 concentrates signal in a handful of features, making the quantization process more signal-preserving and aligned with anisotropic quantization principles. This compatibility also facilitates integration into practical ANN systems (e.g., DiskANN (Jayaram Subramanya et al., 2019)) that already combine graph-based search with PQ, suggesting that CSRv2 can further lower system-level memory while maintaining high recall. However, as discussed in Appendix H.2, while PQ presents an interesting avenue, it necessitates adaptation to function effectively with MRL/CSRv2’s hierarchical representations, which we leave for future work.

### v2

@@ 1 Introduction
In the era of large foundation models, the quality of embeddings has become a decisive factor shaping downstream performance across tasks such as retrieval, classification and recommendation. Yet the dominant practice still relies on dense representations with thousands of dimensions (e.g., 2048 – 8192). While highly expressive, such embeddings incur substantial costs in storage, memory, and inference latency. These inefficiencies are magnified in large-scale and real-time deployments, where embedding computation and serving often dominate system throughput. As models scale further, embedding efficiency emerges as a central bottleneck, which limits both web-scale applications and deployment on resource-constrained platforms such as mobile and edge devices.
@@ 3.1 Tackling Massive Dead Neurons with k-Annealing
(b) Dead neuron during training
(c) Accuracy vs active dim
@@
We find this approach effectively maintains a low dead-neuron rate during training. As shown in Figure 2(b), although dead neurons rise slightly at target sparsity, their final proportion is far lower than training directly with $k_{\mathrm{final}}$. This indicates that a curriculum schedule provides richer gradients and avoids collapse into the dead-neuron regime. Similar to simulated annealing, a larger $k_{\mathrm{init}}$ promotes exploration and diverse activations, while annealing gradually sharpens embeddings toward the ultra-sparse regime. Figure 2(c) confirms this, as k-annealing yields consistent performance gains across sparsity levels.
@@ C.3 Implementation Details
To ensure a fair comparison between MRL and CSRv2 on the MTEB benchmark, particularly with respect to domain alignment, we select a backbone model that is not natively supported by MRL: e5-Mistral-7B (Wang et al., 2023). We then finetune this model on a carefully curated collection of multi-domain datasets. Specifically, the training data is drawn from three complementary sources: (i) datasets included in the MTEB benchmark (Muennighoff et al., 2022), (ii) the embedding training data collection curated by the Sentence Transformers team [^5] https://huggingface.co/datasets/sentence-transformers/embedding-training-data , and (iii) a suite of public retrieval datasets introduced in Zhang et al. (2025a). During preprocessing, we first deduplicate datasets that appear across multiple collections. Subsequently, following the natural supervision strategy outlined in Section 3.2, we sample up to 20,000 sentence pairs per dataset, resulting in a consolidated training corpus of approximately one million examples.
We finetune e5-Mistral-7B on this corpus using a batch size of 2048, which is a scale commonly adopted by existing MRL-compatible models. Full details of the hyperparameter configuration are provided in Table 5. In contrast, the Qwen3-Embedding-4B model (Zhang et al., 2025b) already incorporates native MRL support; thus, no additional finetuning is required for this backbone.
@@ D.1 Evaluation Metrics
Following the methodology established by Kusupati et al. (2022), we adopt 1-nearest neighbor (1-NN) accuracy as the primary metric for evaluating visual representations. This metric is computed using FAISS (Jacob et al., 2018) with exact L2 distance search. In contrast to classification accuracy, which depends on the specific architecture and training procedure of a downstream classifier, 1-NN accuracy provides a direct assessment of whether semantically similar instances are embedded in close proximity within the representation space. Consequently, it serves as a more model-agnostic and training-free probe of intrinsic representation quality.
@@ G.3 Analysis on Unbalanced Weightable Settings
We propose MRL-reweight, where we follow MRL’s settings and set the following weights $\{5,4,3,2,1,\ldots,1\}$ for dimensions 2, 4, 8, 16, 32, and so on, up to 4096. As shown in Table 17, applying larger weights at earlier stages does, to some extent, improve the performance of MRL on low-dimensional scales. However, while MRL-reweight offers some improvements, the performance does not quite reach the level of CSRv2. We hypothesize that this discrepancy arises because sparse vectors, which are more comprehensive in capturing feature combinations, are more difficult to achieve with a truncated representation (e.g., retaining only the first 8 values).
@@ H.3 Potential Applications of CSRv2 in Vector Quantization
Vector quantization (VQ) methods, including Product Quantization (PQ) (Jegou et al., 2010), Optimized Product Quantization (Ge et al., 2013), and more recent anisotropic schemes such as AVQ (Guo et al., 2020), are central to real-world large-scale vector search systems where memory footprint, latency, and hardware efficiency are critical. While our main work focuses on the role of ultra-sparse representations in improving retrieval quality and compute efficiency, it is worth noting that CSRv2 is highly compatible with these widely-used quantization techniques.
CSRv2’s ultra-sparse structure, activating only $k\in\{2,4,8\}$ dimensions out of a large latent space, naturally complements vector quantization methods such as PQ and AVQ. As only a few coordinates are non-zero, quantization can be applied exclusively to these active values (or their indices), enabling a two-stage compression pipeline of sparsity + quantization that substantially reduces both memory and lookup cost. Unlike dense embeddings (e.g., MRL), where quantization error spreads across all dimensions, CSRv2 concentrates signal in a handful of features, making the quantization process more signal-preserving and aligned with anisotropic quantization principles. This compatibility also facilitates integration into practical ANN systems (e.g., DiskANN (Jayaram Subramanya et al., 2019)) that already combine graph-based search with PQ, suggesting that CSRv2 can further lower system-level memory while maintaining high recall. However, as discussed in Appendix H.2, while PQ presents an interesting avenue, it necessitates adaptation to function effectively with MRL/CSRv2’s hierarchical representations, which we leave for future work.